OpenStudy (anonymous):

35,21,11,5,3 whats the next three numbers and why?

OpenStudy (m):

1

OpenStudy (m):

5?

OpenStudy (m):

5, 11, 21? lol

OpenStudy (anonymous):

4,9,18 35-21=14 21-11=10 11-5=6 so An-[A(n-1)-An]-4 will be next term

OpenStudy (anonymous):

I think M was right with 5,11,21. you add -14,-10,-6,-2. So you are adding 4 to what you add each time so next you would add 2,6,10. 3+2=5, 5+6=11, 11+10=21.